Order: NPH insulin 10 units and Humulin insulin 3 units subq

Supply: Vials of NPH insulin 100 units/mL and Humulin 100 units/mL  

  

How many units will you have in your syringe?  
Answer #1

Ordered dose for NPH= 10 Units

Ordered dose for Humulin= 3 units

Available dose of NPH =100 U/ml

For 10 units we have to take,x ml= 10/100= 0.1 ml

Available dose of Humulin= 100 U/ml

So for 3 units , x ml= 3/ 100 = 0.03 ml

So , total ml of medicine to be taken= 0.1+0.03 = 0.13 ml

units to be administered= 13 units

30 units insulin syringe is better for this patient, so can read the markings and take accurate dose when comparing to a large insulin syringe.